OpenStudy (pokemon23):

Solve each equation for x: 4x+9=2x+31 Solve each equation for x: 4x+9=2x+31 @Mathematics

OpenStudy (saifoo.khan):

4x+9=2x+31 x = 11

OpenStudy (pokemon23):

Saifoo again

OpenStudy (saifoo.khan):

:D

OpenStudy (pokemon23):

i subtracted 9 from both sides

OpenStudy (saifoo.khan):

hehe

OpenStudy (lilg132):

4x + 9 = 2x + 31 4x - 2x = 31 - 9 2x = 22 x = 11